OpenStudy (anonymous):

In the website Purple math- Graphing Logarithmic Functions: Examples (page 3 of 3) at the bottom is a graph of y = ln(–x) Isn't y = ln(–x) the same as log base e of (-x) ? I thought that was undefined.

OpenStudy (anonymous):

It is yes. However if x is negative than values can come out.

OpenStudy (anonymous):

the logarithm of a negative integer is not undefined.. it just isn't real.. it's a complex number.. ln(-1) is \[i*\pi\]

OpenStudy (anonymous):

E.g. if x is -2, ln(-x) = ln(-(-2)) = ln(2) = 0.693.. As simonsmiley has pointed out, if you're working with complex numbers, Euler's formula will help you find that ln(-1) = i*pi